if 100cm of hydrogen is sparked with 40cm3 of oxygen to produce H2O at s.t.p, calculate the volume of the gas that is in excess


1
Expert's answer
2022-05-09T10:12:03-0400

2H2 + O2 =2 H2O

n(H2) = 0.1 L / 2*22.4 mol/L = 0.0022 moles

n(O2) = 0.04 L / 22.4 mol/L = 0.0017 moles

H2 - is in excess

V (H2) =( 0.04L • 2•22.4 mol/L) / 22.4 mol/L = 0.08 L